OpenStudy (anonymous):

The equation sinx = √ 2/2 is an identity. True or False?

OpenStudy (aum):

sinx = √ 2/2 is an equation that is good only for some 'x'. An identity will be true for ALL 'x'.

OpenStudy (anonymous):

it was false. But thank you!

OpenStudy (aum):

That is exactly what I said above. sinx = √ 2/2 is an EQUATION from which you should have concluded it is NOT an identity.

OpenStudy (aum):

sin^2(x) + cos^2(x) = 1 is an identity because it is true for ALL x. sinx = √ 2/2 is not an identity because it is true only for some x.